The researcher from Q2 decided that they want to follow up with the 6 children one year after they have completed the intervention program to see how they are now coping with negative emotions. The same method of assessment was used to examine coping responses. The data for the one-year follow-up list is listed below (along with results from Q2). Before One-Year Person Follow-U After Treatment Child Treatment Total Y 297 n=6 k 3 N 18 G 65 17 10 14 4 4 12 SSbefore 11.33

A. Use a repeated measures ANOVA with a=0.05 to determine whether there are any significant mean differences among the three conditions.

B. Calculate n2 to determine the effect size.
Answer #1


Date: 04/05/2019 Answer To test the hypothesis is that there any significant differences among the three treatments means atBefore After Follow-up 4 ty 4 4 3 25 4.17 119 14 26 4.33 134 Sum Average2.33 Sum sq Grand Total Sum sq is 297 Sum is 65 The s142 252 262 (65)2 18 = 249.50-234.72 -14.78 The sum of square totalis. (GT) SST -Sumof squareGT - 2 = 297-(45) = 297-234 .72The degree of freedom for between groups is, =3-1 The degree of freedom error is =18-3 -15 The degree of freedom total is, d,BG dfsG 14.78 7.39 The mean square error is, MSE 47.50 15 3.17 The F-test statistics is, MSBG obt MSE 7.39 3.17 - 2.33The F-test statistics is 2.33 The complete ANOVA table is, Source of VariationSS MS 7.39 2.33 14.78 2 Treatment Error Total 4The F-test statistics is 2.33 is less than F-critical value which is 3.68, the null hypothesis is not rejected at 5% level of